New cars depreciate about 12% of their value each year. Determine the original cost of the car if it depreciated $3,600 in the first year.


1
Expert's answer
2021-01-27T17:20:12-0500

"Solution: \\
let \, the\, price\, of\, the\, car\, be = x \\
depreciation = 12\% \, of\, x\\
depreciation = \$3600 \\
\Rightarrow \frac{12}{100}*x = 3600 \\
\Rightarrow \frac{12x}{100} = 3600 \\
\Rightarrow 12x = 3600 * 100 \\
x = \frac{360000}{12} \\
\Rightarrow x = \$30000 \\

\therefore \, the\, orignal \, cost\, of\, the\, car\,is\, \$30000"